Trimble Lawn mowing bought a new mower for $10,000. The new mower will increase the annual revenues by $50,000, and increase the operating costs by $20,000.   Trimble is using straight-line depreciation over a 5-year economic life. If the tax rate is 30%, please calculate the incremental after-tax cash inflow:
Answer #1

Cash flow net of tax =(50000-20000)*(1-0.3) =21000

Depreciation tax shield =(10000/5)*30% =600

Incremental after-tax CF per year =21000+600 =21600